# Floer-theoretic filtration on Painlevé Hitchin systems

### v1

@@ 3. Construction of actions and filtration computation
In order to define the Painlevé Hitchin systems, one needs to choose further an effective divisor $D$ of length $4$ on $\mathbb{C}P^{1}$. We may and will choose it so that its point of support with the highest multiplicity is $z=0$. The Painlevé Hitchin systems $\mathcal{M}^{PX}$ (where $X$ stands for an element of the set $I,\ldots,VI$) arise as holomorphic symplectic moduli spaces of certain meromorphic Higgs bundles with polar divisor (bounded from above by) $D$, and with fixed polar part at $D$. The holomorphic symplectic form is called Atiyah–Bott $2$-form, and denoted by $\Omega_{I}^{PX}$. Partitions of $4$ and a further choice of the type of the polar part (semi-simple or with non-trivial nilpotent part) at each point of $D$ lead to various possibilities commonly referred to as the Painlevé I, II, III, IV, V and VI cases (with some further labels attached to the cases II, III, IV and V, depending on the authors). We will denote the corresponding Painlevé Hitchin systems by $\mathcal{M}^{PI},\ldots,\mathcal{M}^{PVI}$. Each of them is birational to $\operatorname{Tot}(K)$, and the birational isomorphism identifies the Liouville $2$-form $\Omega_{I}$ with the Atiyah–Bott $2$-form $\Omega_{I}^{PX}$. In addition, they carry fibrations
$h\colon\mathcal{M}^{PX}\to\mathcal{B}^{PX}\cong\mathbb{C}$
turning them into completely integrable systems whose generic fibers are abelian varieties. We construct equivariant $\mathbb{C}^{*}$-actions case by case in Sections 3.13.5. In Section 3.6 we show that in the cases not covered by our constructions, equivariant $\mathbb{C}^{*}$-actions do not exist. In Section 3.7 we treat the complex $2$-dimensional moduli spaces of rank $3,4$ and $6$ parabolic Higgs bundles over $\mathbb{C}P^{1}$ with three logarithmic singularities, corresponding to the root systems $\widetilde{E}_{6},\widetilde{E}_{7}$ and $\widetilde{E}_{8}$ respectively.
According to [KLSS23, Theorem A], the Liouville symplectic form (16) is equal to the Atiyah–Bott holomorphic symplectic form $\Omega_{I}^{PX}$. [^8] Strictly speaking, this was done in the de Rham complex structure rather than the Dolbeault structure needed here; however, a straightforward modification of that proof gives the Dolbeault case too. See also [Mar94, Proposition 7.12] for a related result in the Dolbeault case. It follows immediately from this relationship, that:
@@ 3.1. Painlevé I case
According to [ISS18, Theorem 1.3], the associated Hitchin system $h:\mathcal{M}^{PI}\rightarrow B^{PI}$ has a unique singular fiber, of type $II$ in the Kodaira classification (that is, a genus 00 curve with one cuspidal point, having a local equation analytically equivalent to $y^{2}=x^{3}$). The above local form simplifies to
@@
There exists an $I$-holomorphic $\mathbb{C}^{*}$-action on $\mathcal{M}^{PI}$, equivariant with respect to the action $b\mapsto t^{6}b$ on $\mathcal{B}^{PI}$. It acts on $\Omega_{I}^{PI}$ with weight $5$.
@@
Now, with the above choices we find
@@
Let us determine the $\mathbb{C}^{*}$-fixed locus of $Y:=\mathcal{M}^{PI}$ and their weight decompositions. Due to 2.1,
@@
thus is a cuspidal genus 00 curve. It contains, but is not equal999as the fixed locus is smooth to the $\mathbb{C}^{*}$-fixed locus. Due to the cohomology rank count (combination of (4) and (10)) the fixed locus consists of 2 points,
@@
One of these, say $F_{0}$, has to be the cuspidal point, since $\mathrm{Core}(\mathcal{M}^{PI})\setminus F$ consists of smooth $\mathbb{C}^{*}$-flowlines between $F_{0}$ and $F_{1}$. One of the points is the minimum $F_{\mathrm{min}}$ and the other is the maximum $F_{\mathrm{max}}$ for the moment map $H:Y\rightarrow\mathbb{R}$ of the action $S^{1}\leq\mathbb{C}^{*}$. Now the decomposition
@@
has a negative and a positive weight, where the negative corresponds to the $\mathbb{C}^{*}$-flow from $F_{min}$, and the positive goes outside the core. Due to 2.3, and the last statement of Proposition 3.3, $k_{2}-k_{1}=5,$ in particular $k_{2}\geq 6.$ Due to the first statement of Proposition 3.3, we have $k_{2}|6$, hence $k_{2}=6,$ and so $k_{1}=1.$ Thus, the singular point $F_{0}$ cannot be $F_{\mathrm{max}}$, as otherwise, in local coordinates $(x,y)$ around the $F_{0}=(0,0)$ core would be a (irreducible) cuspidal curve invariant under the $(tx,t^{6}y)$ action, so of type $y=x^{6},$ which is smooth, contradiction. Thus, $F_{1}=F_{\mathrm{max}}$, and the $\mathbb{C}^{*}$-flow within the core is weight-1, hence not torsion (i.e. fixed under a cyclic subgroup of $\mathbb{C}^{*}$). On the other hand, the other $\mathbb{C}^{*}$-flowline emanating from $F_{1}$, of weight $6$, represents the Hitchin section of $h$, consisting of irregular Higgs bundles of the form
@@
on $\mathcal{O}_{\mathbb{C}P^{1}}(-1)\oplus\mathcal{O}_{\mathbb{C}P^{1}}(1)$, using the isomorphism $K_{\mathbb{C}P^{1}}(4\cdot\{0\})\cong\mathcal{O}_{\mathbb{C}P^{1}}(2)$ that maps $z^{-4}\operatorname{d}\!z$ to the degree $2$ polynomial $10\cdot z^{2}+0\cdot z+$ 1. Indeed, as one readily checks, the choice $b=0$ gives rise to a spectral curve that is non-singular and ramified over $z=0$, and has a cusp point over $z=\infty$, and then the Hitchin section intersects it in the ramification point $(z,\zeta)=(0,0)$. The weight decomposition of minimum $F_{0}=F_{\mathrm{min}}$
@@
has only positive weights, and as before, $l_{1}+l_{2}=5.$ Moreover, both of these weight spaces correspond to the $\mathbb{C}^{*}$-flow which goes outside of the core (as due to the equivariant tubular neighbourhood theorem, their image under the exponential map has to be smooth), hence as before $l_{1}|6,\ l_{2}|6.$ Thus, up to a permutation, $l_{1}=2,\ l_{2}=3$.
@@
hence $\mathscr{F}_{1/6}\supset H^{2}(Y)$ and $\mathscr{F}_{1/3}=H^{*}(Y).$ On the other hand, due to 2.5, $1\in H^{0}(F_{\mathrm{min}})$ is not in the filtration until $\lambda=1/3,$ so together with the previous we have:
@@
Their shifts, computed via (15), give the columns of the $E_{1}$-page of the Morse–Bott–Floer spectral sequence, shown on Figure 1. In it, we show all differentials $\operatorname{d}_{r},\ r\geq 1$ stapled together. The column $H^{*}(B_{1})$ is determined by 2.10, as the intersection form of $\mathcal{M}^{PI}$ is trivial. Indeed, this is true as $H_{2}(\mathcal{M}^{PI})=[\mathrm{Core}(Y)],$ and its self-intersection is zero, being a fiber class (of the Hitchin map).
@@
Due to the last statement of Theorem 2.9, the spectral sequence determines the filtration – we see that the class in $H^{2}(\mathcal{M}^{PI})$ gets annihilated by the column with period $1/6$, hence lies in $\mathscr{F}_{1/6}.$ Similarly, $H^{0}$ lies in $\mathscr{F}_{1/3},$ hence we recover Equation 20.
@@ 3.2. Painlevé II case
It follows from [ISS18, Theorem 1.1] that the corresponding Hitchin moduli space $\mathcal{M}^{PII}$ has one singular fiber, of type $III$, namely consisting of two genus 00 components intersecting in a single $A_{3}$ type curve singularity also known as a tacnode, i.e., with a local equation equivalent to $y^{2}=x^{4}$.
@@
The space of quadratic differentials $q$ as in the Proposition will be called irregular Hitchin base in the Painlevé II case, and denoted by $\mathcal{B}^{PII}$. For any $b\in\mathbb{C}$ we define the curve $\Sigma^{PII}_{b}$ by the equation
(21) $\zeta^{2}+\left(-a^{2}z^{-8}+bz^{-4}\right)(\operatorname{d}\!z)^{\otimes 2}=0.$
@@
With notation as in Proposition 3.3 the action on (21) is
(22) $t^{2w(\zeta)}\zeta^{2}-a^{2}t^{-6w(z)}z^{-8}(\operatorname{d}\!z)^{\otimes 2}+b^{\prime}t^{-2w(z)}z^{-4}(\operatorname{d}\!z)^{\otimes 2}.$
To preserve the family $\{\Sigma^{PII}_{b}\}_{b\in\mathbb{C}}$ we need to impose the condition
@@
We then see that in order for (22) to follow from (21) we must have $b^{\prime}=t^{4}b$. ∎
@@ 3.3. Painlevé II (Flaschka–Newell) case
There exists a $I$-holomorphic $\mathbb{C}^{*}$-action on $\mathcal{M}^{PIIFN}$, equivariant with respect to the action $b\mapsto t^{4}b$ on $\mathcal{B}^{PIIFN}$. It acts on $\Omega_{I}^{PIIFN}$ with weight $3$.
@@
Let us determine the $\mathbb{C}^{*}$-fixed locus of $Y:=\mathcal{M}^{PII}$ and their weight decompositions; this does not depend on which of the two cases above we discuss, as in both cases the geometry of $\mathbb{C}^{*}$-action is the same. Given the Hitchin fibration $h:Y\rightarrow\mathcal{B}^{PII}$, the core is (due to 2.1)
@@
As the fixed locus is smooth and its cohomology rank is equal to the cohomology of the total space, we either have 3 fixed points, or a fixed $\mathbb{C}P^{1}$ and a point on the other one. But the latter is impossible, as the two $\mathbb{C}P^{1}$ in the core share the tangent space at their intersection point, which is a weight-0 space, meaning the other $\mathbb{C}P^{1}$ has to be fixed as well, contradiction. Thus, the fixed locus
@@
condition (2.3 and Proposition 3.6), $F_{2}$ then has the same decomposition as $F_{1},$ in particular it has all positive weights, giving another minimum point, which is impossible.101010recall 2.4. Thus, we have
@@
If $H_{2}$ corresponds to flow within $\mathbb{C}P^{1}$’s, then the weight decomposition of $F_{1}$ (and $F_{2}$) would have $H_{-2},$ thus $H_{5}$ weight space as well. The latter would then flow outside of the core, but $5\nmid 4,$ contradiction.111111recall that due to equivariance of $h:Y\rightarrow\mathcal{B}^{PII}$ all outer weights have to divide the weight of the base $\mathcal{B}^{PII},$ which is due to Proposition 3.6 equal to 4. Thus, $H_{1}$ in (25) corresponds to the flow within the $\mathbb{C}P^{1}$’s, hence
(26) $T_{F_{1}}Y=H_{-1}\oplus H_{4},\ \ T_{F_{2}}Y=H_{-1}\oplus H_{4}.$
and $H_{2}$ in (25) goes outside of the core. It follows from [Sza21, Lemma 2 (4)] that one of the two $\mathbb{C}P^{1}$ components of the core, say the one containing $F_{2}$, parameterizes flags at the logarithmic parabolic point $z=\infty$ for a Higgs bundle that has zero residue at that point, corresponding to the choice $b=0$ in the Hitchin base. This component therefore arises from blowing up the cuspidal singular point of the spectral curve lying over $z=\infty$. The $\mathbb{C}^{*}$-flowline of weight $4$ emanating from the other fixed point $F_{1}$ is then the Hitchin section, just as for Painlevé I. Now, computation of the indices (11) gives:121212where, as before we denote $\mu_{\lambda+}:=\mu_{\lambda+\delta}$ for $01\gg\delta>$ 0
@@
hence $\mathscr{F}_{1/4}\supset H^{2}(Y),$ and $\mathscr{F}_{1/2}=H^{*}(Y)$. Due to 2.5 and (25), $1\in H^{0}(F_{\mathrm{min}})=H^{0}(F_{0})$ is not in the filtration until $\lambda=1/2,$ so together with the previous we have:
@@
where one of the components of $Y_{4}$ is the Hitchin section and the second and third component in $Y_{2}$ is $Y_{4},$ as $\mathbb{Z}/2\leq\mathbb{Z}/4$. Thus, their intersections with the energy hypersurfaces (13) are
@@
where each $E_{i}^{2}=-2$, as it is a Lagrangian curve, thus by the Weinstein neighbourhood theorem, its normal bundle is isomorphic to its cotangent bundle, hence the self-intersection is the negative Euler characteristic, i.e. $-2$ ($E_{i}=\mathbb{C}P^{1}$). Intersection $E_{1}\cdot E_{2}=2$ as they are tangential in order 2.

### v2

@@ 3. Construction of $\mathbb{C}^{*}$-actions and filtration computation
In order to define the Painlevé Hitchin systems, one needs to choose further an effective divisor $D$ of length $4$ on $\mathbb{C}P^{1}$. We may and will choose it so that its point of support with the highest multiplicity is $z=0$. The Painlevé Hitchin systems $\mathcal{M}^{PX},X\in\{I,\ldots,VI\}$ arise as holomorphic symplectic moduli spaces of certain meromorphic Higgs bundles with polar divisor (bounded from above by) $D$, and with fixed polar part at $D$. The holomorphic symplectic form is called Atiyah–Bott $2$-form, and denoted by $\Omega_{I}^{PX}$. Partitions of $4$ and a further choice of the type of the polar part (semi-simple or with non-trivial nilpotent part) at each point of $D$ lead to various possibilities commonly referred to as the Painlevé I, II, III, IV, V and VI cases (with some further labels attached to the cases II, III, IV and V, depending on the authors). We will denote the corresponding Painlevé Hitchin systems by $\mathcal{M}^{PI},\ldots,\mathcal{M}^{PVI}$. According to [KLSS23, Theorem A], for each $X$ there is a birational map
$f_{X}\colon\mathcal{M}^{PX}\to\operatorname{Tot}(K)$
such that [^10] Strictly speaking, this was done in the de Rham complex structure rather than the Dolbeault structure needed here; however, a straightforward modification of that proof gives the Dolbeault case too. See also [Mar94, Proposition 7.12] for a related result in the Dolbeault case.
(17) $f_{X}^{*}\Omega_{I}=\Omega_{I}^{PX}.$
In addition, they carry fibrations
(18) $h\colon\mathcal{M}^{PX}\to\mathcal{B}^{PX}\cong\mathbb{C}$
turning them into completely integrable systems whose generic fibers are abelian varieties. We construct equivariant $\mathbb{C}^{*}$-actions case by case in Sections 3.13.5. In Section 3.6 we show that in the cases not covered by our constructions, equivariant $\mathbb{C}^{*}$-actions do not exist. In Section 3.7 we treat the complex $2$-dimensional moduli spaces of rank $3,4$ and $6$ parabolic Higgs bundles over $\mathbb{C}P^{1}$ with three logarithmic singularities, corresponding to the root systems $\widetilde{E}_{6},\widetilde{E}_{7}$ and $\widetilde{E}_{8}$ respectively. From (16) and (17), it follows immediately that:
@@ 3.1. Painlevé I case
According to [ISS18, Theorem 1.3], the associated Hitchin system $h:\mathcal{M}^{PI}\rightarrow B^{PI}$ has a unique singular fiber of type $II$ in the Kodaira classification, namely a genus 00 curve with one cuspidal point, having a local equation analytically equivalent to $y^{2}=x^{3}$. The above local form simplifies to
@@
There exists a holomorphic $\mathbb{C}^{*}$-action on $\mathcal{M}^{PI}$, equivariant with respect to the action $b\mapsto t^{6}b$ on $\mathcal{B}^{PI}$. It acts on $\Omega_{I}^{PI}$ with weight $5$.
@@
Now, with the above choices, we find
@@
Let us determine the $\mathbb{C}^{*}$-fixed loci of $Y:=\mathcal{M}^{PI}$ and their weight decompositions. Due to 2.1,
@@
thus is a cuspidal genus 00 curve. It contains, but is not equal111111as the fixed locus is smooth to the $\mathbb{C}^{*}$-fixed locus. Due to the cohomology rank count (combination of (4) and (10)) the fixed locus consists of 2 points,
@@
One of these, say $F_{0}$, has to be the cuspidal point, since $\mathrm{Core}(Y)\setminus F$ consists of smooth $\mathbb{C}^{*}$-flowlines between $F_{0}$ and $F_{1}$. One of the points is the minimum $F_{\mathrm{min}}$ and the other is the maximum $F_{\mathrm{max}}$ for the restriction of the moment map $H:Y\rightarrow\mathbb{R}$ of the action $S^{1}\leq\mathbb{C}^{*}$ on the core. Now the decomposition
@@
has a negative and a positive weight, where the negative corresponds to the $\mathbb{C}^{*}$-flow from $F_{min}$, and the positive goes outside the core. Due to 2.3, and the last statement of Proposition 3.3, $k_{2}-k_{1}=5,$ in particular $k_{2}\geq 6.$ Due to the first statement of Proposition 3.3, we have $k_{2}|6$, hence $k_{2}=6,$ and so $k_{1}=1.$ Thus, the singular point $F_{0}$ cannot be $F_{\mathrm{max}}$, as otherwise, in local coordinates $(x,y)$ around the $F_{0}=(0,0)$ core would be a (irreducible) cuspidal curve invariant under the $(tx,t^{6}y)$ action, so of type $y=x^{6},$ which is smooth, contradiction. Thus, $F_{1}=F_{\mathrm{max}}$, and the $\mathbb{C}^{*}$-flowline within the core is weight-1, hence not torsion (i.e. fixed under a cyclic subgroup of $\mathbb{C}^{*}$). On the other hand, the $\mathbb{C}^{*}$-flowline emanating from $F_{1}$, of weight $6$, represents the Hitchin section of $h$, consisting of irregular Higgs bundles of the form
@@
on $\mathcal{O}_{\mathbb{C}P^{1}}(-1)\oplus\mathcal{O}_{\mathbb{C}P^{1}}(1)$, using the isomorphism $K_{\mathbb{C}P^{1}}(4\cdot\{0\})\cong\mathcal{O}_{\mathbb{C}P^{1}}(2)$ that maps $z^{-4}\operatorname{d}\!z$ to the degree $2$ polynomial $10\cdot z^{2}+0\cdot z+$ 1. Indeed, as one readily checks, the choice $b=0$ gives rise to a spectral curve that is non-singular and ramified over $z=0$, and has a cusp point over $z=\infty$, and then the Hitchin section intersects it in the ramification point $(z,\zeta)=(0,0)$.
The weight decomposition of minimum $F_{0}=F_{\mathrm{min}}$
@@
has only positive weights, and as before, $l_{1}+l_{2}=5.$ Moreover, both of these weight spaces correspond to the $\mathbb{C}^{*}$-flow which goes outside of the core (as due to the equivariant tubular neighborhood theorem, their image under the exponential map has to be smooth), hence as before $l_{1}|6,\ l_{2}|6.$ Thus, up to a permutation, $l_{1}=2,\ l_{2}=3$.
@@
hence $\mathscr{F}_{1/6}\supset H^{2}(Y)$ and $\mathscr{F}_{1/3}=H^{*}(Y).$ On the other hand, due to 2.5, $1\in H^{0}(F_{\mathrm{min}})$ is not in the filtration until $\lambda=1/3,$ so, together with the previous, we have:
@@
Their shifts, computed via (15), give the columns of the $E_{1}$-page of the Morse–Bott–Floer spectral sequence, shown on Figure 1. In it, we show all the differentials $\operatorname{d}_{r},\ r\geq 1$ stapled together. The column $H^{*}(B_{1})$ is determined by 2.10, as the intersection form of $\mathcal{M}^{PI}$ is trivial. Indeed, this is true as $H_{2}(\mathcal{M}^{PI})=[\mathrm{Core}(Y)],$ and its self-intersection is zero, being a fiber class (of the Hitchin map).
@@
Due to the last statement of Theorem 2.9, the spectral sequence determines the filtration – we see that the class in $H^{2}(\mathcal{M}^{PI})$ gets annihilated by the column with period $1/6$, hence lies in $\mathscr{F}_{1/6}.$ Similarly, $H^{0}$ lies in $\mathscr{F}_{1/3},$ hence we recover the filtration (22) indeed.
@@ 3.2. Painlevé II case
It follows from [ISS18, Theorem 1.1] that the corresponding Hitchin moduli space $\mathcal{M}^{PII}$ has one singular fiber, of type $III$, namely consisting of two genus 00 components intersecting in a single $A_{3}$ type curve singularity also known as a tacnode, with a local equation equivalent to $y^{2}=x^{4}$.
@@
The space of quadratic differentials $q$ as in the Proposition will be called irregular Hitchin base in the Painlevé II case, and denoted by $\mathcal{B}^{PII}$. For any $b\in\mathbb{C}$ we define the spectral curve by the equation
(23) $\Sigma^{PII}_{b}=\{(z,\zeta)\colon\;\zeta^{2}+\left(-a^{2}z^{-8}+bz^{-4}\right)(\operatorname{d}\!z)^{\otimes 2}=0\}.$
@@
With notation as in Proposition 3.3, the action on (23) is
(24) $t^{2w(\zeta)}\zeta^{2}-a^{2}t^{-6w(z)}z^{-8}(\operatorname{d}\!z)^{\otimes 2}+b^{\prime}t^{-2w(z)}z^{-4}(\operatorname{d}\!z)^{\otimes 2}=0$
Thus, to preserve the family $\{\Sigma^{PII}_{b}\}_{b\in\mathbb{C}}$ of spectral curves, we need to impose the condition
@@
We then see that in order for (24) to be the equation of spectral curve  (23), we must have $b^{\prime}=t^{4}b$. ∎
@@ 3.3. Painlevé II (Flaschka–Newell) case
There exists a holomorphic $\mathbb{C}^{*}$-action on $\mathcal{M}^{PIIFN}$, equivariant with respect to the action $b\mapsto t^{4}b$ on $\mathcal{B}^{PIIFN}$. It acts on $\Omega_{I}^{PIIFN}$ with weight $3$.
@@
Let us determine the $\mathbb{C}^{*}$-fixed loci of $Y:=\mathcal{M}^{PII}$ and their weight decompositions; this does not depend on which of the two cases above we discuss, as in both cases the geometry of the $\mathbb{C}^{*}$-action is the same. Given the Hitchin fibration $h:Y\rightarrow\mathcal{B}^{PII}$, the core is (first equality is due to 2.1)
@@
As the fixed locus is smooth and its cohomology rank is equal to the cohomology of the total space, we either have 3 fixed points, or a fixed $\mathbb{C}P^{1}$ and a point on the other one. But the latter is impossible, as the two $\mathbb{C}P^{1}$ in the core share the tangent space at their intersection point, which is a weight00-0- 0 space, meaning the other $\mathbb{C}P^{1}$ would have to be fixed as well, contradiction. Thus, the fixed locus
@@
condition (2.3 and Proposition 3.6), $F_{2}$ then has the same decomposition as $F_{1},$ in particular it has all positive weights, giving another minimum point, which is impossible.121212recall 2.4. Thus, we have
@@
If $H_{2}$ corresponds to flow within $\mathbb{C}P^{1}$’s, then the weight decomposition of $F_{1}$ (and $F_{2}$) would have $H_{-2},$ thus (because of 2.3) $H_{5}$ weight space as well. The latter would then flow outside of the core, but $5\nmid 4,$ contradiction.131313recall that due to equivariance of $h:Y\rightarrow\mathcal{B}^{PII}$ all outer weights have to divide the weight of the base $\mathcal{B}^{PII},$ which is due to Proposition 3.6 equal to 4. Thus, $H_{1}$ in (27) corresponds to the flow within the $\mathbb{C}P^{1}$’s, hence, due to 2.3 we have weight decompositions
(28) $T_{F_{1}}Y=H_{-1}\oplus H_{4},\ \ T_{F_{2}}Y=H_{-1}\oplus H_{4},$
and $H_{2}$ in (27) goes outside of the core.
It follows from [Sza21, Lemma 2 (4)] that one of the two $\mathbb{C}P^{1}$ components of the core, say the one containing $F_{2}$, parametrizes flags at the logarithmic parabolic point $z=\infty$ for a Higgs bundle that has zero residue at that point, corresponding to the choice $b=0$ in the Hitchin base. This component therefore arises from blowing up the cuspidal singular point of the spectral curve lying over $z=\infty$. The $\mathbb{C}^{*}$-flowline of weight $4$ emanating from the other fixed point $F_{1}$ is then the Hitchin section, just as for Painlevé I.
Now, computation of the indices (11) gives:141414where, as before we denote $\mu_{\lambda+}:=\mu_{\lambda+\delta}$ for $01\gg\delta>$ 0
@@
hence $\mathscr{F}_{1/4}\supset H^{2}(Y),$ and $\mathscr{F}_{1/2}=H^{*}(Y)$. Due to 2.5 and (27), $1\in H^{0}(F_{\mathrm{min}})=H^{0}(F_{0})$ is not in the filtration until $\lambda=1/2,$ so, together with the previous, we have:
@@
where one of the components of $Y_{4}$ is the Hitchin section, and the second and third components in $Y_{2}$ is $Y_{4},$ as $\mathbb{Z}/2\leq\mathbb{Z}/4$. Thus, their intersections with the energy hypersurfaces (13) are
@@
where each $E_{i}^{2}=-2$, as it is a Lagrangian curve, thus by the Weinstein neighborhood theorem, its normal bundle is isomorphic to its cotangent bundle, hence the self-intersection is the negative Euler characteristic, i.e. $-2$ ($E_{i}=\mathbb{C}P^{1}$). Intersection $E_{1}\cdot E_{2}=2$ as they are tangential in order 2.